OpenStudy (anonymous):

Find the indicated term of each expansion: Fifth term of (2a+3b)^10

OpenStudy (amistre64):

\[\binom{10}{?} (2a)^{10-5}(3b)^{0+5}\]

OpenStudy (amistre64):

i think thats a 10C5 ....

OpenStudy (amistre64):

10.9.8.7.6 --------- 5.4.3.2 10.9.8.7 --------- 5.4 10.9.2.7 --------- 5 2.9.2.7 = 4(63) = 12 + 240 = 232

OpenStudy (amistre64):

lol, 252

OpenStudy (amistre64):

dbl chk with someone smarter tho ;)